OpenStudy (anonymous):

please help, trying to find center of a circle: (x - 4)^2 + (y + 10)^2 = 121

OpenStudy (acxbox22):

(x−h)^2+(y−k)^2=r^2 is the standard equation of a circle where (h,k) is the center and r is the radius

OpenStudy (acxbox22):

sub in the values from your equation for the values of h and k and you will have your center

OpenStudy (anonymous):

the signs won't flip right? since there already the same as the standard

OpenStudy (anonymous):

Standard equation for Circle is: \[(x-h)^2+(y-k)^2 = r^2\] \((h,k)\) is the center of circle and \(r\) is radius..

OpenStudy (anonymous):

Compare your equation with this standard equation, and find \(h\),\(k\) and \(r\)..

OpenStudy (acxbox22):

no only (y+10)^2 term will switch as their is a + not a -

OpenStudy (anonymous):

okay so is the center (-4, 10) or (-4,-10)

OpenStudy (acxbox22):

there is no - in front of 4 since you can directly sub in 4 for h in (x-h)^2

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

okay my laptop won't load the homepage for openstudy pm or should i post here

OpenStudy (anonymous):

another circle problem: i think i have the correct answer but could someone check to see if i did it correctly? all i have to do is find the center. (x + 3)^2 + y^2 = 45 my guess is (-3,0)

OpenStudy (anonymous):

Great, yes it is right answer.. :) Good.. :)